OpenStudy (anonymous):

The table below shows two equations: Equation 1 |2x – 3| + 5 = 4 Equation 2 |5x + 3| – 10 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution, and equation 2 has solutions x = 2, -3.2. The solutions to equation 1 are x = 1, 2, and equation 2 has no solution. The solutions to equation 1 are x = 1, 2, and equation 2 has solutions x = 2, -3.2.

OpenStudy (amoodarya):

\[|2x – 3| + 5 = 4\\|2x – 3| = 4-5\\|2x – 3| =-1 <0\]

OpenStudy (amoodarya):

\[|somthing|\ge 0\]

OpenStudy (amoodarya):

so eq 1 has no answer

OpenStudy (anonymous):

So its a or b

OpenStudy (amoodarya):

\[|5x + 3| – 10 = 3\\|5x + 3| = 3+10\\|5x + 3|=13\] equation 2 has answer to solve \[|somethimg|=+number\\something =\pm number\\like-this\\|u|=5 \rightarrow u=\pm 5\] can you go on ?

OpenStudy (anonymous):

No, sorry

OpenStudy (anonymous):

Got it, it's B

OpenStudy (amoodarya):

\[|5x+3|=13\\5x+3=\pm13\\first\\5x+3=13 \rightarrow5x=13-3=10 \rightarrow x=\frac{10}{5}=2\\second\\5x+3=-13 \ \rightarrow 5x=-16 \rightarrow x=\frac{-16}{5}\]